OpenStudy (curry):

Naturalists find that the population of some kinds of predatory animals vary periodically, Assume that the population of foxes in a certain forest varies sinusoidally with time. Records started being kept when time t= 0 years. A minimum number, 200 foxes occured when t = 2.9 years. The next maximum, 800 foxes, occured at t = 5.1 year. sketch a graph of this sinusoid write an equation expressing the # of foxes as f(x) of time t predict pop when t=7 (show on graph that this is correct>>Foxes are declared to be an endangered species when their population drops below 200. Between what 2 nonega

OpenStudy (mani_jha):

. Forgive me, I m not good at drawing. The amplitude seems to be A= (800-200)/2=300. Time period(T) = (2.9+1.05)-0=3.95 \[\omega=2\pi/T\] y=Asin(wt)+200. Maybe this is it.